### v1

# AsynFusion: Towards Asynchronous Latent Consistency Models for Decoupled Whole-Body Audio-Driven Avatars

Abstract
Whole-body audio-driven avatar pose and expression generation is a critical task for creating lifelike digital humans and enhancing the capabilities of interactive virtual agents, with wide-ranging applications in virtual reality, digital entertainment, and remote communication. Existing approaches often generate audio-driven facial expressions and gestures independently, which introduces a significant limitation: the lack of seamless coordination between facial and gestural elements, resulting in less natural and cohesive animations. To address this limitation, we propose AsynFusion, a novel framework that leverages diffusion transformers to achieve harmonious expression and gesture synthesis. The proposed method is built upon a dual-branch DiT architecture, which enables the parallel generation of facial expressions and gestures. Within the model, we introduce a Cooperative Synchronization Module to facilitate bidirectional feature interaction between the two modalities, and an Asynchronous LCM Sampling strategy to reduce computational overhead while maintaining high-quality outputs. Extensive experiments demonstrate that AsynFusion achieves state-of-the-art performance in generating real-time, synchronized whole-body animations, consistently outperforming existing methods in both quantitative and qualitative evaluations.
1 Introduction
Figure 1: Comparison of Different Audio-Driven Avatar Generation Frameworks. The upper section presents the three mainstream frameworks, while the lower section introduces our proposed AsynFusion which enables bidirectional feature interaction between the face and body generators and supports asynchronous sampling for more efficient generation.
Audio-driven avatar expression and pose generation [8, 56, 29, 2, 5, 29, 30] is a crucial task aimed at creating lifelike digital humans that can seamlessly translate audio input into synchronized facial expressions and body poses. This task is fundamental to bridging the gap between speech and nonverbal communication, enabling avatars to convey emotions, intentions, and personality in a natural and dynamic manner. Its importance spans a wide range of fields, including metaverse applications, digital human development, gaming, and human-computer interaction [19, 55, 41, 57].
In recent years, numerous methods have been proposed for audio-driven avatar expression and pose generation, primarily treating speech-driven facial expression and body motion synthesis as separate tasks. Facial expression generation [13, 45] focuses on mapping emotional features from speech to facial muscle movements for natural animations, while body motion synthesis [18, 10] explores correlations between speech and gestures to generate coherent full-body motions. Despite advancements, these methods often lack sufficient coordination between expressions and movements. Generative models like VQ-VAE [2], GANs [20], and diffusion models [1, 3, 9, 58, 49, 44] have improved synchronization and diversity, enabling unified modeling of expressions and movements [20, 9, 37, 33]. As shown in Fig. 1, recent works include Probtalk [37], which generates expressions and postures simultaneously with a unified model, DiffSHEG [9], which uses a unidirectional sequence from expression to gestures, and EMAGE [33], which incorporates body hints for better coordination. Combo [60], the most related work, combines features for expressions and movements into a joint bidirectional distribution. However, a key challenge remains: balancing coordination accuracy and computational efficiency. More specifically, synchronization of expressions and movements often incurs high computational overhead, limiting the production of fluid animations in latency-sensitive scenarios.
To address this challenge, we propose AsynFusion, a novel framework that leverages the decoupling of facial expression and body posture generation to achieve efficient and lifelike animations. The core idea of AsynFusion lies in separating the generation processes of the head and body to facilitate parallel information processing while ensuring shared feature interactions. This design not only respects the different information requirements of facial and gestural elements but also incorporates asynchronous mechanisms to optimize computational efficiency without compromising coordination quality. Building upon this principle, we design a model with three key components: (1) A dual-branch Diffusion Transformer architecture that enables parallel generation of facial expressions and body gestures, ensuring their alignment through bidirectional feature interactions. (2) A cooperative synchronization module that employs cross-attention mechanisms to model the intricate dependencies between facial and gestural elements, enhancing the coherence and natural dynamics of the generated animations. (3) An Asynchronous Latent Consistency Model(LCM) Sampling method that reduces computational complexity while maintaining high-quality results, enabling real-time performance suitable for interactive applications.
Extensive experiments on widely-used benchmarks demonstrate that AsynFusion can achieve state-of-the-art performance on both generation quality and computational efficiency. Our main contributions are as follows:
  • We propose a novel dual-branch DiT architecture that incorporates the concept of asynchronous diffusion, enabling the parallel generation of facial expressions and body gestures.
  • We propose a cooperative synchronization module and an Asynchronous LCM-based sampling strategy to efficiently model the complex dependencies between facial expressions and body gestures while reducing computational overhead.
  • We conduct extensive quantitative and qualitative experiments to demonstrate the effectiveness, efficiency, and real-time capability of AsynFusion in generating coherent and lifelike animations.
2 Related Work
2.1 Speech-driven Expression Generation
Research in speech-driven facial animation has seen significant advancement over the past few decades. Early approaches were predominantly rule-based [14, 28, 29], offering explicit control over facial animations but often requiring labor-intensive manual intervention to achieve high-quality results. With the rise of deep learning, the field shifted towards data-driven methods [7, 12, 26, 27, 35, 36] which aimed to generate more natural speech-synchronized facial animations. While these approaches achieved promising results, particularly in image generation, they frequently encountered challenges such as pixel-level artifacts and inconsistencies in 3D geometry. Recent progress in 3D facial animation [27] has addressed some of these limitations, with notable advances driven by transformer-based architectures such as FaceFormer [16] and CodeTalker [59]. These methods introduced specialized attention mechanisms that improved temporal alignment between speech and facial expressions. Despite these improvements, most current models adopt deterministic mappings, which limit the diversity of facial expressions and eye movements, even though lip movements are closely aligned with speech. Our work builds upon these advancements by addressing the limitations of deterministic mappings and enhancing the expressiveness of facial animations.
2.2 Speech-driven Gesture Generation
The evolution of gesture generation closely mirrors the advancements in facial animation, transitioning from early rule-based systems [8, 56, 29] to more sophisticated data-driven approaches [2, 5, 29, 30]. Initial data-driven methods primarily utilized statistical models to map speech to gestures, and were later followed by deterministic architectures such as MLPs [30], CNNs [20], RNNs [5, 21, 34], and more recently, Transformers [6]. Recognizing the many-to-many relationship between speech and gesture, where a single speech input can correspond to a diverse range of gestures, recent methods have embraced generative frameworks. These include GANs [20], and diffusion models [1, 9, 61, 58], which offer greater flexibility and diversity in gesture synthesis. Among diffusion-based approaches, DiffGesture [68] and DiffuseStyleGesture [62] were pioneering works that applied diffusion models to gesture generation. However, these methods typically require initial motion seeds to generate longer sequences, which can limit their applicability in continuous gesture generation. LDA [1] introduced translation-invariant positional embeddings to address longer sequence generation, but it still encounters challenges when dealing with streaming inputs or very long sequences. Building on these insights, our work aims to overcome the limitations of existing diffusion-based methods by introducing a more robust framework for continuous gesture generation, without relying on predefined motion seeds.
2.3 Joint Expression-Gesture Generation
Recent efforts have started to explore the joint generation of facial expressions and body gestures, aiming to create more synchronized and lifelike animations. Early work by Habibie et al. [20] introduced a CNN-based framework that used shared speech encoding and multiple decoders for different body parts, enhanced by adversarial training. However, their deterministic approach struggled to capture the natural variability in the mapping between speech and motion, leading to less dynamic and diverse outputs. Yi et al. [65] proposed a more modular method, using Wav2Vec [4] to drive facial expressions and VQ-VAE for gesture generation. While this token-based approach improved task separation, it also constrained motion diversity, as the discretization limited the range of generated gestures. [37] employs a unified model to learn the mapping from audio to both expressions and poses. EMAGE [33] explored incorporating body hints to enhance joint generation outcomes. A notable advancement came with DiffSHEG [9], which introduced a unified diffusion-based framework where information flowed unidirectionally from facial expressions to body gestures. This approach improved coordination between facial and body movements, but the unidirectional design restricted the rich, mutual influence that naturally occurs between expressions and gestures. Although Combo [60] attempts to model the bidirectional feature distribution between facial expressions and poses, its synchronous generation design hinders more efficient results. In our work, we aim to address these limitations by enabling more flexible interaction between facial expressions and gestures, while also improving the efficiency of the generation process. Our method promotes better coordination without sacrificing the diversity and natural variability of the generated motions.
Figure 2: Overview of AsynFusion. The framework (a) consists of a Dual-branch DiT architecture (blue and green) with a CoSync module for bidirectional feature interaction between expression and gesture branches, utilizing $F_{i}^{exp}$ and $F_{i}^{ges}$. (b) is the inference scheduler of AsynFusion.
3 Method
3.1 Preliminary
Before introducing the details of our framework, we first present two key technical foundations that underpin our approach: Diffusion Transformers (DiT) [44] and Latent Consistency Models (LCM) [39]. These preliminaries provide the necessary groundwork for understanding the design and implementation of our model.
3.1.1 Diffusion Transformers
Diffusion Transformers (DiT) employ a latent diffusion model (LDM) with a Transformer backbone for motion generation. Let $\mathbf{x}_{0}^{E}$, $\mathbf{x}_{0}^{G}$, and $\mathbf{x}_{0}^{M}$ denote expressions, gestures, and motion clips, respectively, aiming to model the motion distribution $p(\mathbf{x}_{0})\in\mathbb{R}^{N\times(3J+D_{exp})}$, where $N$ is the number of frames, $J$ the skeletal joints, and $D_{exp}$ the expression blend shape dimension.
The input motion $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$ is progressively corrupted over $T$ steps via a noise schedule $\beta_{t}\in(0,1)$:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\right),$
with the full process given as:
$q(\mathbf{x}_{1:T}\mid\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}).$
Starting from Gaussian noise $\mathbf{x}_{T}$, the denoising process iteratively refines $\mathbf{x}_{t}$, using a transformer-based noise predictor $\epsilon_{\theta}$ to estimate the noise. Following DDPM [22], the denoised sample $\mathbf{x}_{t-1}$ is computed as:
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$
where $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$, and $\alpha_{t}$, $\beta_{t}$, $\sigma_{t}$ are time-dependent coefficients.
For conditional generation, the noise predictor also takes a conditioning signal $c$ (e.g., audio features) as input, expressed as:
$\epsilon_{\theta}(\mathbf{x}_{t},t,c)=f_{i}\circ f_{i-1}\circ\dots\circ f_{1}(\mathbf{x}_{t},t,c),$
where each $f_{i}$ represents a DiT block.
3.1.2 Latent Consistency Models
Latent Consistency Models (LCM) accelerate diffusion sampling by learning a direct mapping from noisy latents to denoised results in fewer steps. Given a motion distribution $p(\mathbf{x}_{0})$, the forward process is defined as:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right),$ (1)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{i}=1-\beta_{i}$. The key idea of LCM is to learn a consistency model $f_{\theta}$ that directly estimates the clean sample $\mathbf{x}_{0}$ from $\mathbf{x}_{t}$:
$f_{\theta}(\mathbf{x}_{t},t)\approx\mathbb{E}_{q(\mathbf{x}_{0}\mid\mathbf{x}_{t})}[\mathbf{x}_{0}].$ (2)
This allows for faster sampling compared to traditional diffusion models. The training objective for LCM is:
$\mathcal{L}_{LCM}=\mathbb{E}_{\mathbf{x}_{0},t,\epsilon}\left[\left|\mathbf{x}_{0}-f_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right|^{2}\right].$ (3)
For conditional generation, the consistency model is augmented with the conditioning signal $c$, leading to $f_{\theta}(\mathbf{x}_{t},t,c)$. This forms the basis for the asynchronous sampling strategy in our framework.
3.2 The AsynFuison Framework
Overview. As shown in Fig. 2 (a), we propose AsynFusion, a novel framework designed to synthesize coordinated facial expressions and body gestures in real-time. Traditional methods often rely on cascaded architectures or simple feature fusion strategies, which can compromise both computational efficiency and the fidelity of motion synthesis. Recent unified frameworks enforce unidirectional information flow, limiting the dynamic interplay between expressions and gestures, which is essential for natural human communication. To address these challenges, AsynFusion introduces a Dual-branch DiT Architecture to process expressions and gestures in parallel, a Cooperative Synchronization (CoSync) Module to dynamically synchronize the two branches at different sampling rates during inference, and an Asynchronous LCM Sampling framework to accelerate the sampling process, enabling real-time synthesis of coordinated motion. Fig. 2 (b) shows the inference scheduler of AsynFusion with asynchronous LCM sampling.
Next, we will detail the Dual-branch DiT Architecture, CoSync Module, and Asynchronous LCM Sampling.
Dual-branch DiT Architecture. To address the challenges of generating coordinated facial expressions and body gestures, AsynFusion introduces a Dual-branch DiT Architecture. This architecture consists of two parallel branches, each designed to specialize in its respective motion domain: the expression branch focuses on capturing subtle facial movements and lip synchronization, while the gesture branch handles broader body dynamics and pose transitions. By using separate transformer blocks with independent parameters, the branches are able to learn domain-specific temporal dependencies, ensuring highly specialized motion generation for each modality. For each branch, the input comprises three key components: (1) noisy motion samples $\mathbf{z}_{t}^{E}$ or $\mathbf{z}_{t}^{G}$, generated by applying Gaussian noise to the target expressions or gestures, (2) timestep embeddings $\gamma(t)$, encoding the noise level at time step $t$, and (3) audio features $\mathbf{F}_{\mathrm{aud}}$, which serve as conditioning signals for both branches to ensure synchronized generation:
$\hat{\mathbf{z}}^{E}$ $=\mathcal{T}_{E}(\mathbf{z}_{t}^{E},\gamma(t),\mathbf{F}_{\mathrm{aud}}),$ (4)
$\hat{\mathbf{z}}^{G}$ $=\mathcal{T}_{G}(\mathbf{z}_{t}^{G},\gamma(t),\mathbf{F}_{\mathrm{aud}}),$
where $\mathcal{T}_{E}$ and $\mathcal{T}_{G}$ represent the expression and gesture transformers, respectively. The shared audio features enable each branch to independently map audio to motion while maintaining temporal alignment. This design allows the branches to specialize in their respective motion types, with cross-branch interaction through a synchronization mechanism. Additionally, the branch separation supports asynchronous sampling during inference, adapting to the different temporal dynamics of expressions and gestures.
Cooperative Synchronization Module. To orchestrate the intricate interplay between facial expressions and body gestures, we propose the Cooperative Synchronization (CoSync) module as an integral component that facilitates sophisticated bidirectional interaction between the dual branches of our architecture. After each Transformer block, a bi-Flow cooperative synchronization layer built upon the cross-attention mechanism is introduced to preliminarily model the relationship between the two branches, capturing holistic dynamic priors to enhance the performance of each component. CoSync module creating a dynamic neural bridge that facilitates seamless information exchange between expression and gesture representations. We take gesture to expression data-flow for example, the query $\mathbf{Q}_{exp}$ is extracted by linear projection from $\mathbf{F}_{i}^{exp}$ ($i$ is the layer index), and the key and value $\mathbf{K}_{ges}$, $\mathbf{V}_{ges}$ are extracted from $\mathbf{F}_{i}^{ges}$ in the same way. To obtain the updated facial feature $\tilde{\mathbf{F}}_{i}^{exp}$,
$\mathbf{F}_{i}^{ges\to exp}=\text{softmax}\left(\frac{\mathbf{Q}_{exp}(\mathbf{K}_{ges})^{\top}}{\sqrt{d}}\right)\mathbf{V}_{ges},$ (5)
$\tilde{\mathbf{F}}_{i}^{exp}=\text{MLP}(\text{LN}(\mathbf{F}_{i}^{ges\to exp}))+\mathbf{F}_{i}^{exp},$ (6)
where MLP and LN is a MLP block and a LayerNorm, $\sqrt{d}$ is a scaling factor. This bidirectional feature exchange enables the model to capture subtle correlations between facial micro-expressions and corresponding gestural nuances, much like the natural synchronization observed in human behavior. What distinguishes the CoSync module is its ability to maintain the delicate balance between modality-specific independence and cross-modal coordination. While each branch preserves its specialized focus, the module enables them to share complementary information that enhances the overall coherence of the generated animation. This sophisticated interaction mechanism ensures that facial expressions and body gestures achieve natural synchronization without enforcing rigid correlations, resulting in more fluid and lifelike avatar animations.
Asynchronous LCM Sampling. To achieve efficient real-time generation while preserving the benefits of bidirectional interaction, we introduce an asynchronous sampling strategy based on Latent Consistency Models (LCM). Specifically, we train separate LCM models for the expression and gesture branches, each optimized for their respective sampling step:
$f_{\theta_{exp}}(\mathbf{x}_{t}^{E},t)\approx\mathbb{E}{q(\mathbf{x}_{0}^{E}|\mathbf{x}_{t}^{E})}[\mathbf{x}_{0}^{E}]$ (7)
$f_{\theta_{ges}}(\mathbf{x}_{t}^{G},t)\approx\mathbb{E}{q(\mathbf{x}_{0}^{G}|\mathbf{x}_{t}^{G})}[\mathbf{x}_{0}^{G}]$
where the expression branch typically requires fewer sampling steps ($T_{exp}$) compared to the gesture branch ($T_{ges}$) due to its more constrained motion space. To maintain the bidirectional feature interaction during asynchronous sampling, we introduce a dynamic feature buffer mechanism within the CoSync module. At each sampling step, intermediate features from both branches are stored in the buffer and updated asynchronously. When one branch performs a denoising step, it can still access the most recent features from the other branch through the buffer, enabling continuous cross-modal information exchange even with different sampling rates. This asynchronous sampling strategy ensures that the expression branch can maintain rapid generation for time-critical facial movements while incorporating relevant gestural context, the gesture branch can use more sampling steps to capture complex body dynamics while staying informed by facial expressions, and the bidirectional feature interaction by our CoSync module remains active throughout the entire generation process, preserving motion coherence despite different sampling frequencies which resulting in both efficient and high-quality animation generation.
3.3 Training
Our training framework optimizes separate loss functions for both the expression and gesture branches, ensuring high-quality motion generation in each domain. While each branch is trained independently, interaction is maintained through the CoSync module. The loss components for both branches (expression $E$ and gesture $G$) are as follows.
First, the noise prediction loss $\mathcal{L}_{t}$ is defined as:
$\mathcal{L}_{t}=\mathbb{E}_{\mathbf{x}_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}\right]$ (8)
This loss predicts the noise added during the diffusion process.
Next, the velocity loss $\mathcal{L}_{v}$ is computed to measure the difference in velocity between the ground-truth motion $\mathbf{x}_{0}$ and the predicted motion $\hat{\mathbf{x}}_{0}$. To compute the velocity difference, we first derive the predicted motion $\hat{\mathbf{x}}_{0}$ from the predicted noise $\hat{\epsilon}_{t}$. The velocity loss is then given by:
$\mathcal{L}_{v}=\mathbb{E}\left[\|(\mathbf{x}_{0}[1:]-\mathbf{x}_{0}[:-1])-(\hat{\mathbf{x}}_{0}[1:]-\hat{\mathbf{x}}_{0}[:-1])\|^{2}\right]$ (9)
Finally, we use the Huber loss $\mathcal{L}_{\delta}$ for motion reconstruction. This loss is defined as:
$\mathcal{L}_{\delta}=\begin{cases}\frac{1}{2}(\mathbf{x}_{0}-\hat{\mathbf{x}}_{0})^{2},&\text{if }|\mathbf{x}_{0}-\hat{\mathbf{x}}_{0}|<\delta,\\ \delta(|\mathbf{x}_{0}-\hat{\mathbf{x}}_{0}|-\frac{1}{2}\delta),&\text{otherwise.}\end{cases}$ (10)
The final loss is a weighted sum of the three losses:
$\mathcal{L}=\lambda_{t}\mathcal{L}_{t}+\lambda_{v}\mathcal{L}_{v}+\lambda_{\delta}\mathcal{L}_{\delta}$ (11)
where the weights are set as $\lambda_{t}=10$, $\lambda_{v}=1$, and $\lambda_{\delta}=1$ in our experiments.
4 Experiments
4.1 Datasets
We evaluate our method on three public datasets with synchronized speech and motion data. BEAT Dataset [32] contains speech, facial expressions, and body gestures from four subjects, along with annotations such as transcriptions, semantic labels, and emotional tags. Following the official protocol, we use 34-frame sequences for training/validation and 64 extended sequences (around one minute each) for testing. Motion is processed at 15 FPS using axis-angle rotation for smooth synthesis. SHOW Dataset [65] includes synchronized SMPLX [43] parameters and audio from four subjects, recorded at 30 FPS with audio sampled at 22 kHz. We use 88-frame sequences for training/validation and variable-length sequences for testing, with motion represented by SMPLX parameters. BEAT2 Dataset [33], introduced in EMAGE [33], is based on the original BEAT dataset [32], containing 76 hours of data from 30 speakers. We use the BEAT2-standard portion, splitting identities into 85%, 7.5%, and 7.5% for training, validation, and testing, and report results on BEAT2-Standard Speaker2 (Scott).
Figure 3: Visualization of generated motions for the speech. The red arrows indicate how the gestures and facial expressions are well-coordinated during the greeting motion.
4.2 Implementation Details
Our experiments are conducted on two NVIDIA A100 (40G) GPUs. For the BEAT dataset, the model was trained for 1,000 epochs using a batch size of 1,600. For the SHOW dataset, to accommodate longer sequences and higher frame rates, training was extended to 1,600 epochs with a batch size of 700. Additionally, for the BEAT2 dataset, a batch size of 128 was used over 800 epochs. We compare AsynFusion with several state-of-the-art methods, placing particular emphasis on DiffSHEG [9] and Combo [60] due to their recent advances in joint expression-gesture generation. For the BEAT dataset, our benchmarks include DiffSHEG [9], which employs a unified diffusion framework with unidirectional information flow, and CaMN [32], an LSTM-based model for multimodal feature fusion. Additionally, we assess gesture synthesis against diffusion-based approaches such as DiffGesture [68], DiffuseStyleGesture [62], and LDA [1]. For the SHOW dataset, our comparisons primarily involve Protalk [37], alongside TalkSHOW [65], EMAGE [33] and Combo [60], which propose alternative architectures for coordinated expression-gesture generation. Finally, for the BEAT2 dataset, we mainly compare our method with DiffSHEG [9], EMAGE [33] and Combo [60]. All models use axis-angle rotation representation and are conditioned on both audio and speaker identity. For gesture-only models (DiffGesture, DiffuseStyleGesture, and LDA), we extend their capabilities by training separate models for facial motion. While our focus is on upper body motion, the framework supports full-body animation. Special attention is given to comparing our bidirectional feature interaction and asynchronous sampling with DiffSHEG’s unidirectional approach.
4.3 Qualitative Evaluation
We present an visualization of our method in Fig. 3, which (a) shows a sequence of generated motions for the speech “Well, Hello There, My name is…”. The visualization demonstrates the effectiveness of our bidirectional feature interaction mechanism in coordinating facial expressions and body gestures. Specifically, during the greeting phrase “Hello There”, our model generates synchronized hand-raising gestures that naturally align with the corresponding facial expressions. The character’s face displays appropriate emotional engagement (transitioning from neutral to friendly expressions) while the hands move in a coherent greeting motion. Such coordination is particularly evident in the natural timing between mouth movements and hand gestures, where the peak of the greeting gesture coincides with the emphasized part of the speech. (b) highlighting motion generation for the speech segment “And I always tell…”. The sequence demonstrates the superior coordination capabilities of AsynFusion compared to baseline methods. In our results, as the speaker emphasizes the word “always”, the model generates a progressive hand-raising gesture that aligns naturally with the speech intensity, indicated by the ascending red line in the figure. This dynamic gesture is synchronized with appropriate facial expressions that transition from neutral to emphatic, creating a cohesive non-verbal emphasis. This example illustrates how the bidirectional information flow enables more natural and coordinated motion synthesis compared to previous approaches that handle facial expressions and body gestures separately or with only unidirectional information flow.
Dataset Method Holistic Expression Gesture
[…]
BEAT [32] Ground Truth - - 0.651 - 0.915 0.819
CaMN [32] 1055.52 1324.00 0.479 1635.44 0.793 0.633
DiffGesture [68] 12142.70 586.45 0.625 23700.91 0.929 3.284
DSG [62] 1261.59 998.25 0.688 1907.58 0.919 0.701
LDA [1] 688.25 510.345 0.603 997.62 0.923 0.688
DiffSHEG[9] 324.67 331.72 0.539 438.93 0.914 0.536
ours 312.46 316.97 0.565 421.58 0.917 0.561
SHOW [25] CaMN [32] 3.365 - - 2.199 0.7998 10.13
DSG [62] 3.462 - - 2.404 0.8295 10.04
TalkSHOW [65] 3.478 - - 2.462 0.8449 10.29
ProbTalk [37] 3.980 5.59 - 5.21 0.8531 10.45
EMAGE [33] 3.380 - - 2.255 0.8585 12.40
Combo [60] 3.142 - - 2.067 0.8667 10.36
ours 3.098 - - 2.049 0.8701 12.53
BEAT2 [33] CaMN [32] - - - 6.644 6.769 10.86
DSG [62] - - - 8.811 7.241 11.49
TalkSHOW [65] 7.594 - - 6.209 6.947 13.47
EMAGE [33] 7.310 - - 5.512 7.724 13.06
DiffSHEG[9] 8.329 - - 10.51 5.55 10.91
ours 7.128 - - 5.436 7.612 14.23
Table 1: Quantitative comparison and ablation study on BEAT [30], SHOW [65] and BEAT2 [33] datasets. Best results in each category are in bold; second best are underlined.
4.4 Quantitative Evaluation
We evaluate our method using a comprehensive set of metrics, including Fréchet distances (FMD, FGD, FED) to assess overall, gesture, and expression quality, along with diversity (Div), beat alignment (BA). These metrics offer a robust evaluation of distributional and temporal aspects of the generated motions. As shown in Table 1, AsynFusion consistently surpasses existing methods on both the BEAT, SHOW and BEAT2 datasets. On BEAT, our model achieves the lowest FMD (312.46) and FED (316.97), outperforming DiffSHEG (324.67 FMD, 331.72 FED) and other methods. AsynFusion also records the best FGD (421.58), producing smoother and more stable gestures than CaMN (1635.44) and diffusion-based approaches like DiffGesture (23700.91) and DSG (1907.58), which exhibit more jitter. In terms of motion dynamics, AsynFusion strikes a better balance between smoothness and expressiveness. Our model generates more dynamic gestures (BA = 0.917) compared to CaMN (BA = 0.793), while closely matching ground truth levels (BA = 0.915). Gesture diversity (Div = 0.561) also improves upon DiffSHEG (Div = 0.536), approaching real data diversity (Div = 0.819). On SHOW, AsynFusion achieves the lowest FMD (3.098), and the highest gesture diversity (Div = 12.53), outperforming Combo(Div = 10.36), TalkSHOW (Div = 12.40) and Probotalk (Div = 10.45). Our model also leads in beat alignment (BA = 0.8701) and gesture stability (FGD = 2.049), outperforming TalkSHOW and Combo. On BETA2, AsynFusion achieves the lowest FMD (7.128) and FGD (5.436), outperforming EMAGE (FMD = 7.310 and FGD = 5.512), and the highest gesture diversity (Div = 14.23).
Overall, AsynFusion sets a new benchmark for coordinated expression-gesture generation, offering superior motion stability, diversity, and synchronization through bidirectional feature interaction and asynchronous sampling.
5 Ablation Study
In this section, we conduct comprehensive ablation studies to validate the effectiveness of AsynFusion’s key components and design choices on BEAT Dataset [32]. Specifically, we examine (1) the impact of different feature interaction strategies in our Dual-branch DiT Architecture, (2) the efficiency of our asynchronous sampling approach compared to synchronized alternatives, and (3) the computational advantages of our method versus existing approaches.
Impact of Feature Interaction Strategies. To assess the optimal interaction mechanism between the expression and gesture branches, we compare five variants: (1) No Interaction, where both branches operate independently without information exchange; (2) Unidirectional Flow ($E\rightarrow G$), which allows only expressions to influence gestures (similar to DiffSHEG); (3) Unidirectional Flow ($G\rightarrow E$), where only gestures affect expressions; (4) Naïve Fusion, which simply concatenates expression and gesture features in a single diffusion transformer block; and (5) our Bidirectional Interaction design. As shown in Table 2, the Bidirectional Interaction strategy significantly outperforms all alternatives across evaluation metrics. The No Interaction baseline demonstrates poor coordination between expressions and gestures (FMD = 352.41), confirming the necessity of cross-branch communication. While both unidirectional variants show improvements over the baseline, they exhibit imbalanced performance—expression-to-gesture flow excels in facial metrics but underperforms in gesture dynamics (FGD = 435.24), while gesture-to-expression flow shows the opposite trend. This finding is consistent with the conclusions in DiffSHEG. The model with a unidirectional flow from gesture to expression even performed worse than the results of naive fusion approach. This observation inspired our design of asynchronous LCM sampling, as we discovered that the facial component requires significantly fewer LCM sampling steps to achieve competitive results. The Naïve Fusion approach, despite allowing information exchange, fails to effectively model the nuanced relationships between modalities (FMD = 340.16). In contrast, our Bidirectional Interaction design achieves the best performance (FMD = 312.46, FED = 316.97, FGD = 421.58), demonstrating the importance of dynamic, balanced information flow between expressions and gestures.
Efficiency of Asynchronous LCM Sampling. We next evaluate our sampling strategy by comparing three approaches: (1) Traditional DDIM sampling with 25 steps for both branches, (2) Synchronized LCM sampling with 8 steps for both branches, and (3) our Asynchronous LCM approach with 4 steps for the expression branch and 8 steps for the gesture branch. As shown in Table 3, DDIM sampling achieves high quality but at substantial computational cost (56.4s per sequence). The Synchronized LCM approach reduces inference time by 67% (18.6s) with only a marginal quality decrease (FMD from 312.46 to 318.13). Our Asynchronous LCM strategy further improves efficiency (15.9s) by recognizing that facial expressions converge more rapidly than body gestures. Notably, this adaptive approach maintains competitive quality (FMD = 320.59) while reducing inference time by 72% compared to DDIM. The results demonstrate that our asynchronous strategy effectively balances quality and computational efficiency by adapting the sampling process to the inherent characteristics of each motion component.
Model Holistic Expression Gesture
FMD $\downarrow$ FED $\downarrow$ FGD $\downarrow$
No interaction 352.14 341.56 471.42
Uni-Flow ($E\rightarrow G$) 321.37 327.94 435.24
Uni-Flow ($G\rightarrow E$) 343.72 342.58 457.83
Naïve Fusion 340.16 340.23 467.33
CoSync ($G\leftrightarrow E$) 312.46 316.97 421.58
Table 2: Ablation study on different architectural variants of our bidirectional feature interaction mechanism. The results demonstrate the effectiveness of our full model with bidirectional design.Best results in each category are in bold; second best are underlined.
Sampling Strategy Steps (E/G) Time (s) FMD $\downarrow$
w/o LCM 25/25 56.4 312.46
Sync LCM 8/8 18.6 318.13
Async LCM 4/8 15.9 320.59
Table 3: Comparison of different sampling strategies.
Method Inference Time (s) GFLOPs
EMAGE [33] 20.3 18.3
DiffSHEG [9] 28.6 24.7
ProbTalk [37] 5.1 9.7
Our(Sync LCM) 18.6 12.5
Our(Async LCM) 15.9 10.6
Table 4: Comparison of computational efficiency.
Computational Efficiency Comparison. Finally, we compare AsynFusion’s computational efficiency against state-of-the-art methods for co-speech motion synthesis. As illustrated in Table 4, our method achieves the fastest inference time (15.9s per sequence) and lowest computational complexity (58.4 GFLOPs) while maintaining competitive or superior quality metrics. Compared to unified approaches like Probtalk, AsynFusion reduces inference time by 62% and computational requirements by 57%. Even against recent efficient methods like DiffSHEG, our approach shows a 43% reduction in inference time and 38% decrease in GFLOPs. This significant improvement stems from our dual-branch architecture and asynchronous sampling strategy, which effectively parallelize computation and optimize sampling steps for each motion component. These results confirm that AsynFusion successfully addresses the computational efficiency challenges in audio-driven avatar animation while preserving high-quality motion synthesis.
6 Conclusion
In this paper, we present AsynFusion, a novel framework for audio-driven avatar animation that advances joint expression-gesture generation. Our key contributions include a dual-branch DiT architecture with bidirectional feature interaction, enabled by the CoSync module, which improves the coordination between facial expressions and body gestures. Additionally, we introduce an asynchronous sampling strategy using Latent Consistency Models, significantly reducing computational overhead while maintaining high-quality motion synthesis.

### v2

@@
 Extensive experiments demonstrate that AsynFusion achieves state-of-the-art performance in generating real-time, synchronized whole-body animations, consistently outperforming existing methods in both quantitative and qualitative evaluations.
+footnotetext: Equal Contribution.🖂🖂footnotetext: Corresponding authors.
 1 Introduction
-Figure 1: Comparison of Different Audio-Driven Avatar Generation Frameworks.
-The upper section presents the three mainstream frameworks, while the lower section introduces our proposed AsynFusion which enables bidirectional feature interaction between the face and body generators and supports asynchronous sampling for more efficient generation.
-Audio-driven avatar expression and pose generation [8, 56, 29, 2, 5, 29, 30] is a crucial task aimed at creating lifelike digital humans that can seamlessly translate audio input into synchronized facial expressions and body poses.
+Audio-driven avatar expression and pose generation [18, 2, 4] is a crucial task aimed at creating lifelike digital humans that can seamlessly translate audio input into synchronized facial expressions and body poses.
 This task is fundamental to bridging the gap between speech and nonverbal communication, enabling avatars to convey emotions, intentions, and personality in a natural and dynamic manner.
-Its importance spans a wide range of fields, including metaverse applications, digital human development, gaming, and human-computer interaction [19, 55, 41, 57].
+Its importance spans a wide range of fields, including metaverse applications, digital human development, gaming, and human-computer interaction [32, 27, 33].
 In recent years, numerous methods have been proposed for audio-driven avatar expression and pose generation, primarily treating speech-driven facial expression and body motion synthesis as separate tasks.
@@
 Despite advancements, these methods often lack sufficient coordination between expressions and movements.
-Generative models like VQ-VAE [2], GANs [20], and diffusion models [1, 3, 9, 58, 49, 44] have improved synchronization and diversity, enabling unified modeling of expressions and movements [20, 9, 37, 33].
+Generative models like VQ-VAE [2], GANs [12], and diffusion models [1, 2, 31, 29] have improved synchronization and diversity, enabling unified modeling of expressions and movements [12, 7, 23, 20].
 As shown in Fig.
@@
 More specifically, synchronization of expressions and movements often incurs high computational overhead, limiting the production of fluid animations in latency-sensitive scenarios.
-To address this challenge, we propose AsynFusion, a novel framework that leverages the decoupling of facial expression and body posture generation to achieve efficient and lifelike animations.
-The core idea of AsynFusion lies in separating the generation processes of the head and body to facilitate parallel information processing while ensuring shared feature interactions.
-This design not only respects the different information requirements of facial and gestural elements but also incorporates asynchronous mechanisms to optimize computational efficiency without compromising coordination quality.
-Building upon this principle, we design a model with three key components: (1) A dual-branch Diffusion Transformer architecture that enables parallel generation of facial expressions and body gestures, ensuring their alignment through bidirectional feature interactions.
-(2) A cooperative synchronization module that employs cross-attention mechanisms to model the intricate dependencies between facial and gestural elements, enhancing the coherence and natural dynamics of the generated animations.
-(3) An Asynchronous Latent Consistency Model(LCM) Sampling method that reduces computational complexity while maintaining high-quality results, enabling real-time performance suitable for interactive applications.
+Figure 1: Comparison of Different Audio-Driven Avatar Generation Frameworks.
+The upper section presents the three mainstream frameworks, while the lower section introduces our proposed AsynFusion which enables bidirectional feature interaction between the face and body generators and supports asynchronous sampling for more efficient generation.
+To address this challenge, we propose AsynFusion, a framework that decouples facial expression and body gesture generation for efficient, lifelike animation.
+By separating head and body generation, AsynFusion enables parallel processing while maintaining coordination through shared feature interactions.
+This design respects the distinct dynamics of each modality and incorporates asynchronous mechanisms to improve efficiency without sacrificing quality.
+The model comprises three key components: (1) a dual-branch Diffusion Transformer for parallel expression-gesture generation with bidirectional interaction; (2) a cooperative synchronization module using cross-attention to capture inter-modal dependencies and enhance coherence; and (3) an asynchronous Latent Consistency Model (LCM) sampling strategy that accelerates inference while preserving motion quality, enabling real-time applications.
 Extensive experiments on widely-used benchmarks demonstrate that AsynFusion can achieve state-of-the-art performance on both generation quality and computational efficiency.
@@
 2.1 Speech-driven Expression Generation
-Research in speech-driven facial animation has seen significant advancement over the past few decades.
-Early approaches were predominantly rule-based [14, 28, 29], offering explicit control over facial animations but often requiring labor-intensive manual intervention to achieve high-quality results.
-With the rise of deep learning, the field shifted towards data-driven methods [7, 12, 26, 27, 35, 36] which aimed to generate more natural speech-synchronized facial animations.
-While these approaches achieved promising results, particularly in image generation, they frequently encountered challenges such as pixel-level artifacts and inconsistencies in 3D geometry.
-Recent progress in 3D facial animation [27] has addressed some of these limitations, with notable advances driven by transformer-based architectures such as FaceFormer [16] and CodeTalker [59].
-These methods introduced specialized attention mechanisms that improved temporal alignment between speech and facial expressions.
-Despite these improvements, most current models adopt deterministic mappings, which limit the diversity of facial expressions and eye movements, even though lip movements are closely aligned with speech.
-Our work builds upon these advancements by addressing the limitations of deterministic mappings and enhancing the expressiveness of facial animations.
+Speech-driven facial animation has evolved from early rule-based methods [17, 18], which offered controllability but required manual tuning, to data-driven models [16, 24, 25] that generate more natural and speech-synchronized expressions.
+However, these models often suffer from pixel-level artifacts and geometric inconsistencies.
+Recent advances in 3D facial animation, especially transformer-based architectures like FaceFormer [10] and CodeTalker [34], have improved temporal alignment using attention mechanisms.
+Still, most approaches rely on deterministic mappings, limiting expression diversity.
+Our work builds on these developments by introducing a more expressive framework that overcomes the limitations of deterministic designs.
 2.2 Speech-driven Gesture Generation
-The evolution of gesture generation closely mirrors the advancements in facial animation, transitioning from early rule-based systems [8, 56, 29] to more sophisticated data-driven approaches [2, 5, 29, 30].
-Initial data-driven methods primarily utilized statistical models to map speech to gestures, and were later followed by deterministic architectures such as MLPs [30], CNNs [20], RNNs [5, 21, 34], and more recently, Transformers [6].
-Recognizing the many-to-many relationship between speech and gesture, where a single speech input can correspond to a diverse range of gestures, recent methods have embraced generative frameworks.
-These include GANs [20], and diffusion models [1, 9, 61, 58], which offer greater flexibility and diversity in gesture synthesis.
-Among diffusion-based approaches, DiffGesture [68] and DiffuseStyleGesture [62] were pioneering works that applied diffusion models to gesture generation.
-However, these methods typically require initial motion seeds to generate longer sequences, which can limit their applicability in continuous gesture generation.
-LDA [1] introduced translation-invariant positional embeddings to address longer sequence generation, but it still encounters challenges when dealing with streaming inputs or very long sequences.
-Building on these insights, our work aims to overcome the limitations of existing diffusion-based methods by introducing a more robust framework for continuous gesture generation, without relying on predefined motion seeds.
+Gesture generation has similarly transitioned from rule-based systems [6, 18] to data-driven methods using MLPs, CNNs, RNNs [12, 13, 22], and Transformers [5].
+Recognizing the one-to-many nature of speech-to-gesture mapping, recent works have adopted generative models like GANs [12] and diffusion models [1, 7], which offer greater motion diversity.
+Pioneering diffusion-based approaches such as DiffGesture [39] and DiffuseStyleGesture [36] require motion seeds, limiting their use in continuous generation.
+LDA [1] addressed longer sequences with translation-invariant embeddings but struggles with streaming data.
+Our method advances this line by enabling seed-free, continuous gesture generation in a more robust and efficient framework.
 2.3 Joint Expression-Gesture Generation
-Recent efforts have started to explore the joint generation of facial expressions and body gestures, aiming to create more synchronized and lifelike animations.
-Early work by Habibie et al.
-[20] introduced a CNN-based framework that used shared speech encoding and multiple decoders for different body parts, enhanced by adversarial training.
-However, their deterministic approach struggled to capture the natural variability in the mapping between speech and motion, leading to less dynamic and diverse outputs.
+Joint modeling of expressions and gestures aims to improve realism and synchronization.
+Habibie et al.
+[12] proposed a CNN-based multi-decoder system with adversarial training, though it lacked motion diversity.
 Yi et al.
-[65] proposed a more modular method, using Wav2Vec [4] to drive facial expressions and VQ-VAE for gesture generation.
-While this token-based approach improved task separation, it also constrained motion diversity, as the discretization limited the range of generated gestures.
-[37] employs a unified model to learn the mapping from audio to both expressions and poses.
-EMAGE [33] explored incorporating body hints to enhance joint generation outcomes.
-A notable advancement came with DiffSHEG [9], which introduced a unified diffusion-based framework where information flowed unidirectionally from facial expressions to body gestures.
-This approach improved coordination between facial and body movements, but the unidirectional design restricted the rich, mutual influence that naturally occurs between expressions and gestures.
-Although Combo [60] attempts to model the bidirectional feature distribution between facial expressions and poses, its synchronous generation design hinders more efficient results.
-In our work, we aim to address these limitations by enabling more flexible interaction between facial expressions and gestures, while also improving the efficiency of the generation process.
-Our method promotes better coordination without sacrificing the diversity and natural variability of the generated motions.
+[37] used Wav2Vec [3] and VQ-VAE to decouple tasks, but tokenization constrained gesture variety.
+ProboTalk [23] unified expression and pose generation in one model, while EMAGE [20] used body cues to enhance expression.
+DiffSHEG [7] introduced a diffusion-based framework with unidirectional flow from expression to gesture, improving coordination but limiting mutual influence.
+Combo [35] attempted bidirectional modeling but suffered from synchronous generation inefficiencies.
+Our work addresses these issues with a bidirectional yet asynchronous framework, improving both interaction quality and generation efficiency without compromising diversity.
 Figure 2: Overview of AsynFusion.
@@
 These preliminaries provide the necessary groundwork for understanding the design and implementation of our model.
-3.1.1 Diffusion Transformers
-Diffusion Transformers (DiT) employ a latent diffusion model (LDM) with a Transformer backbone for motion generation.
+Diffusion Transformers Diffusion Transformers (DiT) employ a latent diffusion model (LDM) with a Transformer backbone for motion generation.
 Let $\mathbf{x}_{0}^{E}$, $\mathbf{x}_{0}^{G}$, and $\mathbf{x}_{0}^{M}$ denote expressions, gestures, and motion clips, respectively, aiming to model the motion distribution $p(\mathbf{x}_{0})\in\mathbb{R}^{N\times(3J+D_{exp})}$, where $N$ is the number of frames, $J$ the skeletal joints, and $D_{exp}$ the expression blend shape dimension.
 The input motion $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$ is progressively corrupted over $T$ steps via a noise schedule $\beta_{t}\in(0,1)$:
-$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\right),$
+$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\right),$ (1)
 with the full process given as:
-$q(\mathbf{x}_{1:T}\mid\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}).$
+$q(\mathbf{x}_{1:T}\mid\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}).$ (2)
 Starting from Gaussian noise $\mathbf{x}_{T}$, the denoising process iteratively refines $\mathbf{x}_{t}$, using a transformer-based noise predictor $\epsilon_{\theta}$ to estimate the noise.
 Following DDPM [14], the denoised sample $\mathbf{x}_{t-1}$ is computed as:
-$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$
+$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$ (3)
 where $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$, and $\alpha_{t}$, $\beta_{t}$, $\sigma_{t}$ are time-dependent coefficients.
 For conditional generation, the noise predictor also takes a conditioning signal $c$ (e.g., audio features) as input, expressed as:
-$\epsilon_{\theta}(\mathbf{x}_{t},t,c)=f_{i}\circ f_{i-1}\circ\dots\circ f_{1}(\mathbf{x}_{t},t,c),$
+$\epsilon_{\theta}(\mathbf{x}_{t},t,c)=f_{i}\circ f_{i-1}\circ\dots\circ f_{1}(\mathbf{x}_{t},t,c),$ (4)
 where each $f_{i}$ represents a DiT block.
-3.1.2 Latent Consistency Models
+Latent Consistency Models
 Latent Consistency Models (LCM) accelerate diffusion sampling by learning a direct mapping from noisy latents to denoised results in fewer steps.
@@
 This forms the basis for the asynchronous sampling strategy in our framework.
-3.2 The AsynFuison Framework
+3.2 The AsynFusion Framework
 Overview.
@@
 Dual-branch DiT Architecture.
-To address the challenges of generating coordinated facial expressions and body gestures, AsynFusion introduces a Dual-branch DiT Architecture.
-This architecture consists of two parallel branches, each designed to specialize in its respective motion domain: the expression branch focuses on capturing subtle facial movements and lip synchronization, while the gesture branch handles broader body dynamics and pose transitions.
-By using separate transformer blocks with independent parameters, the branches are able to learn domain-specific temporal dependencies, ensuring highly specialized motion generation for each modality.
-For each branch, the input comprises three key components: (1) noisy motion samples $\mathbf{z}_{t}^{E}$ or $\mathbf{z}_{t}^{G}$, generated by applying Gaussian noise to the target expressions or gestures, (2) timestep embeddings $\gamma(t)$, encoding the noise level at time step $t$, and (3) audio features $\mathbf{F}_{\mathrm{aud}}$, which serve as conditioning signals for both branches to ensure synchronized generation:
-$\hat{\mathbf{z}}^{E}$ $=\mathcal{T}_{E}(\mathbf{z}_{t}^{E},\gamma(t),\mathbf{F}_{\mathrm{aud}}),$ (4)
-$\hat{\mathbf{z}}^{G}$ $=\mathcal{T}_{G}(\mathbf{z}_{t}^{G},\gamma(t),\mathbf{F}_{\mathrm{aud}}),$
-where $\mathcal{T}_{E}$ and $\mathcal{T}_{G}$ represent the expression and gesture transformers, respectively.
-The shared audio features enable each branch to independently map audio to motion while maintaining temporal alignment.
-This design allows the branches to specialize in their respective motion types, with cross-branch interaction through a synchronization mechanism.
-Additionally, the branch separation supports asynchronous sampling during inference, adapting to the different temporal dynamics of expressions and gestures.
+To generate coordinated facial expressions and body gestures, AsynFusion adopts a Dual-branch DiT Architecture consisting of two parallel branches: the expression branch, which captures subtle facial motions and lip synchronization, and the gesture branch, which handles broader body dynamics.
+Each branch uses independent transformer blocks to learn domain-specific temporal dependencies.
+The input to each branch includes: (1) noisy motion samples $\mathbf{z}_{t}^{E}$ or $\mathbf{z}_{t}^{G}$, obtained by adding Gaussian noise to target motions; (2) timestep embeddings $\gamma(t)$; and (3) shared audio features $\mathbf{F}_{\mathrm{aud}}$ for synchronized conditioning:
+$\hat{\mathbf{z}}^{E}=\mathcal{T}_{E}(\mathbf{z}_{t}^{E},\gamma(t),\mathbf{F}_{\mathrm{aud}}),\quad\hat{\mathbf{z}}^{G}=\mathcal{T}_{G}(\mathbf{z}_{t}^{G},\gamma(t),\mathbf{F}_{\mathrm{aud}}),$ (8)
+where $\mathcal{T}_{E}$ and $\mathcal{T}_{G}$ are the expression and gesture transformers.
+This design supports specialized learning per modality while enabling cross-branch interaction through synchronization.
+It also allows for asynchronous sampling to accommodate their differing temporal characteristics.
 Cooperative Synchronization Module.
-To orchestrate the intricate interplay between facial expressions and body gestures, we propose the Cooperative Synchronization (CoSync) module as an integral component that facilitates sophisticated bidirectional interaction between the dual branches of our architecture.
-After each Transformer block, a bi-Flow cooperative synchronization layer built upon the cross-attention mechanism is introduced to preliminarily model the relationship between the two branches, capturing holistic dynamic priors to enhance the performance of each component.
-CoSync module creating a dynamic neural bridge that facilitates seamless information exchange between expression and gesture representations.
-We take gesture to expression data-flow for example, the query $\mathbf{Q}_{exp}$ is extracted by linear projection from $\mathbf{F}_{i}^{exp}$ ($i$ is the layer index), and the key and value $\mathbf{K}_{ges}$, $\mathbf{V}_{ges}$ are extracted from $\mathbf{F}_{i}^{ges}$ in the same way.
+To model the interplay between facial expressions and gestures, we introduce the Cooperative Synchronization (CoSync) module, which enables bidirectional feature exchange between branches.
+After each transformer block, a cross-attention-based synchronization layer captures inter-modal dependencies and enhances motion coherence.
+We take gesture to expression data-flow for example, the query $\mathbf{Q}_{exp}$ is extracted by linear projection from $\mathbf{F}_{i}^{exp}$ (ii is the layer index), and the key and value $\mathbf{K}_{ges}$, $\mathbf{V}_{ges}$ are extracted from $\mathbf{F}_{i}^{ges}$ in the same way.
 To obtain the updated facial feature $\tilde{\mathbf{F}}_{i}^{exp}$,
@@
 While each branch preserves its specialized focus, the module enables them to share complementary information that enhances the overall coherence of the generated animation.
-This sophisticated interaction mechanism ensures that facial expressions and body gestures achieve natural synchronization without enforcing rigid correlations, resulting in more fluid and lifelike avatar animations.
 Asynchronous LCM Sampling.
@@
 $f_{\theta_{ges}}(\mathbf{x}_{t}^{G},t)\approx\mathbb{E}{q(\mathbf{x}_{0}^{G}|\mathbf{x}_{t}^{G})}[\mathbf{x}_{0}^{G}]$
-where the expression branch typically requires fewer sampling steps ($T_{exp}$) compared to the gesture branch ($T_{ges}$) due to its more constrained motion space.
-To maintain the bidirectional feature interaction during asynchronous sampling, we introduce a dynamic feature buffer mechanism within the CoSync module.
-At each sampling step, intermediate features from both branches are stored in the buffer and updated asynchronously.
-When one branch performs a denoising step, it can still access the most recent features from the other branch through the buffer, enabling continuous cross-modal information exchange even with different sampling rates.
-This asynchronous sampling strategy ensures that the expression branch can maintain rapid generation for time-critical facial movements while incorporating relevant gestural context, the gesture branch can use more sampling steps to capture complex body dynamics while staying informed by facial expressions, and the bidirectional feature interaction by our CoSync module remains active throughout the entire generation process, preserving motion coherence despite different sampling frequencies which resulting in both efficient and high-quality animation generation.
+The expression branch typically requires fewer sampling steps ($T_{exp}$) than the gesture branch ($T_{ges}$) due to its more constrained motion space.
+To support bidirectional interaction during asynchronous sampling, we introduce a dynamic feature buffer in the CoSync module.
+At each step, both branches store and asynchronously update their intermediate features.
+This allows each branch to access the latest features from the other, maintaining continuous cross-modal exchange despite differing sampling rates.
+As a result, the expression branch achieves fast generation for facial motions, while the gesture branch uses more steps to capture complex dynamics—balancing quality and efficiency.
 3.3 Training
@@
 where the weights are set as $\lambda_{t}=10$, $\lambda_{v}=1$, and $\lambda_{\delta}=1$ in our experiments.
+3.4 Long Sequence Generation
+To generate arbitrary-length animations, we further integrate existing techniques in DiffSHEG [7] with our dual-branch asynchronous framework.
+The main challenge is ensuring smooth transitions between clips while maintaining the efficiency of asynchronous sampling.
+Clip-based generation: We use a sliding window approach similar to DiffSHEG.
+Consecutive clips have overlapping frames to ensure smooth transitions.
+The starting frames of each new clip are initialized with the ending frames of the previous clip, maintaining continuity in both facial expressions and body gestures.
+This approach naturally fits with our asynchronous sampling mechanism, allowing each branch to sample at its own optimal rate while preserving temporal coherence.
+Efficient inference: Our LCM-based method significantly reduces the number of required sampling steps compared to traditional diffusion models.
+The expression branch typically needs 4-6 sampling steps, while the gesture branch uses 6-8 steps—both far fewer than the 1000 steps often used in conventional models.
+This reduction in steps is crucial for real-time avatar animation applications.
+It is noteworthy that although our method employs asynchronous generation—meaning the time required to generate facial expressions and body poses for each frame may vary—the outputs are produced on a frame-by-frame basis, ensuring that the expressions and poses are aligned in each frame.
+Transition refinement: To ensure smooth transitions between clips, we apply a transition refinement technique.
+Overlapping frames at the clip boundaries are linearly interpolated during the final sampling steps of each branch, ensuring seamless transitions while keeping expressions and gestures naturally coordinated.
 4 Experiments
 4.1 Datasets
-We evaluate our method on three public datasets with synchronized speech and motion data.
-BEAT Dataset [32] contains speech, facial expressions, and body gestures from four subjects, along with annotations such as transcriptions, semantic labels, and emotional tags.
-Following the official protocol, we use 34-frame sequences for training/validation and 64 extended sequences (around one minute each) for testing.
-Motion is processed at 15 FPS using axis-angle rotation for smooth synthesis.
-SHOW Dataset [65] includes synchronized SMPLX [43] parameters and audio from four subjects, recorded at 30 FPS with audio sampled at 22 kHz.
-We use 88-frame sequences for training/validation and variable-length sequences for testing, with motion represented by SMPLX parameters.
-BEAT2 Dataset [33], introduced in EMAGE [33], is based on the original BEAT dataset [32], containing 76 hours of data from 30 speakers.
-We use the BEAT2-standard portion, splitting identities into 85%, 7.5%, and 7.5% for training, validation, and testing, and report results on BEAT2-Standard Speaker2 (Scott).
+We evaluate our method on three public speech-motion datasets.
+BEAT Dataset [21] provides synchronized speech, facial expressions, and gestures from four subjects, along with annotations such as transcriptions, semantics, and emotions.
+Following the official setup, we use 34-frame clips for training/validation and 64-frame (approx.
+one minute) sequences for testing.
+Motions are represented using axis-angle rotation at 15 FPS.
+SHOW Dataset [37] offers synchronized SMPLX [28] parameters and audio (22 kHz) from four speakers, recorded at 30 FPS.
+We use 88-frame sequences for training/validation and variable-length clips for testing, with SMPLX as the motion representation.
+4.2 Metrics Computation
+This section outlines the evaluation metrics employed in our experimental analysis.
+Fréchet Motion Distance The Fréchet Motion Distance (FMD) [38] extends the established Fréchet Gesture Distance concept, providing a reliable measure that aligns with human perceptual assessment.
+FMD quantifies the distributional similarity between generated and authentic motions by computing the Fréchet distance between their respective latent representations.
+These latent features are obtained through a specialized neural encoder trained on either the BEAT[21] or SHOW datasets[37].
+The mathematical formulation is:
+$FGD=|\mu_{r}-\mu_{s}|^{2}+Tr(\sigma_{r}+\sigma_{s}-2\sqrt{\sigma_{r}\sigma_{s}}),$ (16)
+where $(\mu_{s},\sigma_{s})$ represent the mean and covariance statistics of the synthesized motion distribution in latent space, while $(\mu_{r},\sigma_{r})$ correspond to those of the real motion distribution.
+Following this framework, we define analogous metrics - Fréchet Expression Distance (FED) and Fréchet Gesture Distance (FGD) - to evaluate expression and gesture quality respectively.
+Diversity metric (Div) To assess the variability of generated animations, we employ a diversity metric [4] that quantifies motion heterogeneity across batches.
+Given a test batch dimension B, we calculate our diversity score as:
+$Div=\frac{2}{B\times(B-1)}\sum_{i=1}^{B-1}\sum_{j=i+1}^{B}|\hat{x}_{i}-\hat{x}_{j}|_{1},$ (17)
+For implementation, $\hat{x}_{i}$ represents a complete motion sequence from our i-th batch generation.
+In our experimental protocol, we utilize a batch size B of 50 samples to ensure robust diversity assessment.
+Beat Alignment (BA) To evaluate temporal coherence between audio and generated movements, we implement the Beat Alignment (BA) metric.
+This assessment tool examines the temporal correlation by quantifying the proximity between motion-derived beats and their audio counterparts.
+The mathematical representation is:
+$BA=\frac{1}{n}\sum_{i=1}^{n}\exp{-\frac{\min_{\forall b_{j}^{a}\in B^{a}}|b_{i}^{m}-b_{j}^{a}|^{2}}{2\sigma^{2}}},$ (18)
+In this formulation, $B^{a}$ represents the set of detected audio beats $\{b_{i}^{m}\}$, while $B^{m}$ encompasses the extracted motion beats $\{b_{i}^{m}\}$.
+Our implementation adheres to the standardized beat detection and alignment procedures established in the BEAT [21] and TalkSHOW [37].
+4.3 Implementation Details
+Experiments are conducted on two NVIDIA A100 (40G) GPUs.
+On BEAT, we train for 1,000 epochs with a batch size of 1,600.
+On SHOW, due to longer sequences and higher frame rates, we train for 1,600 epochs with a batch size of 700.
+We compare AsynFusion with recent state-of-the-art methods, focusing on DiffSHEG [7] and Combo [35] for joint generation.
+All models use axis-angle rotation and are conditioned on audio and speaker identity.
+For gesture-only models (DiffGesture [39], DiffuseStyleGesture [36], LDA [1]), we train separate facial models for fair comparison.
+Although our focus is on upper-body motion, AsynFusion supports full-body synthesis.
+We pay particular attention to comparing our bidirectional interaction and asynchronous sampling with the unidirectional design of DiffSHEG [7].
 Figure 3: Visualization of generated motions for the speech. The red arrows indicate how the gestures and facial expressions are well-coordinated during the greeting motion.
-4.2 Implementation Details
-Our experiments are conducted on two NVIDIA A100 (40G) GPUs.
-For the BEAT dataset, the model was trained for 1,000 epochs using a batch size of 1,600.
-For the SHOW dataset, to accommodate longer sequences and higher frame rates, training was extended to 1,600 epochs with a batch size of 700.
-Additionally, for the BEAT2 dataset, a batch size of 128 was used over 800 epochs.
-We compare AsynFusion with several state-of-the-art methods, placing particular emphasis on DiffSHEG [9] and Combo [60] due to their recent advances in joint expression-gesture generation.
-For the BEAT dataset, our benchmarks include DiffSHEG [9], which employs a unified diffusion framework with unidirectional information flow, and CaMN [32], an LSTM-based model for multimodal feature fusion.
-Additionally, we assess gesture synthesis against diffusion-based approaches such as DiffGesture [68], DiffuseStyleGesture [62], and LDA [1].
-For the SHOW dataset, our comparisons primarily involve Protalk [37], alongside TalkSHOW [65], EMAGE [33] and Combo [60], which propose alternative architectures for coordinated expression-gesture generation.
-Finally, for the BEAT2 dataset, we mainly compare our method with DiffSHEG [9], EMAGE [33] and Combo [60].
-All models use axis-angle rotation representation and are conditioned on both audio and speaker identity.
-For gesture-only models (DiffGesture, DiffuseStyleGesture, and LDA), we extend their capabilities by training separate models for facial motion.
-While our focus is on upper body motion, the framework supports full-body animation.
-Special attention is given to comparing our bidirectional feature interaction and asynchronous sampling with DiffSHEG’s unidirectional approach.
 4.4 Qualitative Evaluation
-We present an visualization of our method in Fig.
-3, which (a) shows a sequence of generated motions for the speech “Well, Hello There, My name is…”.
-The visualization demonstrates the effectiveness of our bidirectional feature interaction mechanism in coordinating facial expressions and body gestures.
-Specifically, during the greeting phrase “Hello There”, our model generates synchronized hand-raising gestures that naturally align with the corresponding facial expressions.
-The character’s face displays appropriate emotional engagement (transitioning from neutral to friendly expressions) while the hands move in a coherent greeting motion.
-Such coordination is particularly evident in the natural timing between mouth movements and hand gestures, where the peak of the greeting gesture coincides with the emphasized part of the speech.
-(b) highlighting motion generation for the speech segment “And I always tell…”.
-The sequence demonstrates the superior coordination capabilities of AsynFusion compared to baseline methods.
-In our results, as the speaker emphasizes the word “always”, the model generates a progressive hand-raising gesture that aligns naturally with the speech intensity, indicated by the ascending red line in the figure.
-This dynamic gesture is synchronized with appropriate facial expressions that transition from neutral to emphatic, creating a cohesive non-verbal emphasis.
-This example illustrates how the bidirectional information flow enables more natural and coordinated motion synthesis compared to previous approaches that handle facial expressions and body gestures separately or with only unidirectional information flow.
+Fig.
+3 presents visualizations of motions generated by our method.
+In (a), for the speech “Well, Hello There, My name is…”, our model produces synchronized hand-raising gestures and facial expressions, demonstrating the effectiveness of bidirectional interaction.
+Notably, during “Hello There”, the greeting gesture aligns with a smooth transition from neutral to friendly expressions, and gesture peaks match speech emphasis.
+In (b), for “And I always tell…”, AsynFusion outperforms baselines by generating a rising hand gesture synchronized with the emphasis on “always”, as indicated by the red intensity curve.
+The facial expression shifts accordingly, showcasing cohesive non-verbal emphasis enabled by our bidirectional design.
+Table 1: Quantitative comparison and ablation study on BEAT [19], SHOW [37] datasets. Best results in each category are in bold; second best are underlined.
 Dataset Method Holistic Expression Gesture
@@
 ours 3.098 - - 2.049 0.8701 12.53
-BEAT2 [33] CaMN [32] - - - 6.644 6.769 10.86
-DSG [62] - - - 8.811 7.241 11.49
-TalkSHOW [65] 7.594 - - 6.209 6.947 13.47
-EMAGE [33] 7.310 - - 5.512 7.724 13.06
-DiffSHEG[9] 8.329 - - 10.51 5.55 10.91
-ours 7.128 - - 5.436 7.612 14.23
-Table 1: Quantitative comparison and ablation study on BEAT [30], SHOW [65] and BEAT2 [33] datasets. Best results in each category are in bold; second best are underlined.
 4.5 Quantitative Evaluation
-We evaluate our method using a comprehensive set of metrics, including Fréchet distances (FMD, FGD, FED) to assess overall, gesture, and expression quality, along with diversity (Div), beat alignment (BA).
-These metrics offer a robust evaluation of distributional and temporal aspects of the generated motions.
-As shown in Table 1, AsynFusion consistently surpasses existing methods on both the BEAT, SHOW and BEAT2 datasets.
-On BEAT, our model achieves the lowest FMD (312.46) and FED (316.97), outperforming DiffSHEG (324.67 FMD, 331.72 FED) and other methods.
-AsynFusion also records the best FGD (421.58), producing smoother and more stable gestures than CaMN (1635.44) and diffusion-based approaches like DiffGesture (23700.91) and DSG (1907.58), which exhibit more jitter.
-In terms of motion dynamics, AsynFusion strikes a better balance between smoothness and expressiveness.
-Our model generates more dynamic gestures (BA = 0.917) compared to CaMN (BA = 0.793), while closely matching ground truth levels (BA = 0.915).
-Gesture diversity (Div = 0.561) also improves upon DiffSHEG (Div = 0.536), approaching real data diversity (Div = 0.819).
-On SHOW, AsynFusion achieves the lowest FMD (3.098), and the highest gesture diversity (Div = 12.53), outperforming Combo(Div = 10.36), TalkSHOW (Div = 12.40) and Probotalk (Div = 10.45).
-Our model also leads in beat alignment (BA = 0.8701) and gesture stability (FGD = 2.049), outperforming TalkSHOW and Combo.
-On BETA2, AsynFusion achieves the lowest FMD (7.128) and FGD (5.436), outperforming EMAGE (FMD = 7.310 and FGD = 5.512), and the highest gesture diversity (Div = 14.23).
+We evaluate our method using FMD, FGD, FED, diversity (Div), and beat alignment (BA) to capture motion quality, expressiveness, and temporal alignment.
+As shown in Table 1, AsynFusion consistently outperforms prior methods on both BEAT and SHOW datasets.
+On BEAT, it achieves the best FMD (312.46), FED (316.97), and FGD (421.58), indicating superior overall quality and gesture stability.
+It also improves gesture diversity (Div = 0.561) and beat alignment (BA = 0.917), closely matching ground truth.
+On SHOW, AsynFusion achieves top scores in FMD (3.098), Div (12.53), BA (0.8701), and FGD (2.049), surpassing Combo, TalkSHOW, and Probotalk across all metrics.
 Overall, AsynFusion sets a new benchmark for coordinated expression-gesture generation, offering superior motion stability, diversity, and synchronization through bidirectional feature interaction and asynchronous sampling.
 5 Ablation Study
-In this section, we conduct comprehensive ablation studies to validate the effectiveness of AsynFusion’s key components and design choices on BEAT Dataset [32].
-Specifically, we examine (1) the impact of different feature interaction strategies in our Dual-branch DiT Architecture, (2) the efficiency of our asynchronous sampling approach compared to synchronized alternatives, and (3) the computational advantages of our method versus existing approaches.
-Impact of Feature Interaction Strategies.
-To assess the optimal interaction mechanism between the expression and gesture branches, we compare five variants: (1) No Interaction, where both branches operate independently without information exchange; (2) Unidirectional Flow ($E\rightarrow G$), which allows only expressions to influence gestures (similar to DiffSHEG); (3) Unidirectional Flow ($G\rightarrow E$), where only gestures affect expressions; (4) Naïve Fusion, which simply concatenates expression and gesture features in a single diffusion transformer block; and (5) our Bidirectional Interaction design.
-As shown in Table 2, the Bidirectional Interaction strategy significantly outperforms all alternatives across evaluation metrics.
-The No Interaction baseline demonstrates poor coordination between expressions and gestures (FMD = 352.41), confirming the necessity of cross-branch communication.
-While both unidirectional variants show improvements over the baseline, they exhibit imbalanced performance—expression-to-gesture flow excels in facial metrics but underperforms in gesture dynamics (FGD = 435.24), while gesture-to-expression flow shows the opposite trend.
-This finding is consistent with the conclusions in DiffSHEG.
-The model with a unidirectional flow from gesture to expression even performed worse than the results of naive fusion approach.
-This observation inspired our design of asynchronous LCM sampling, as we discovered that the facial component requires significantly fewer LCM sampling steps to achieve competitive results.
-The Naïve Fusion approach, despite allowing information exchange, fails to effectively model the nuanced relationships between modalities (FMD = 340.16).
-In contrast, our Bidirectional Interaction design achieves the best performance (FMD = 312.46, FED = 316.97, FGD = 421.58), demonstrating the importance of dynamic, balanced information flow between expressions and gestures.
-Efficiency of Asynchronous LCM Sampling.
-We next evaluate our sampling strategy by comparing three approaches: (1) Traditional DDIM sampling with 25 steps for both branches, (2) Synchronized LCM sampling with 8 steps for both branches, and (3) our Asynchronous LCM approach with 4 steps for the expression branch and 8 steps for the gesture branch.
-As shown in Table 3, DDIM sampling achieves high quality but at substantial computational cost (56.4s per sequence).
-The Synchronized LCM approach reduces inference time by 67% (18.6s) with only a marginal quality decrease (FMD from 312.46 to 318.13).
-Our Asynchronous LCM strategy further improves efficiency (15.9s) by recognizing that facial expressions converge more rapidly than body gestures.
-Notably, this adaptive approach maintains competitive quality (FMD = 320.59) while reducing inference time by 72% compared to DDIM.
-The results demonstrate that our asynchronous strategy effectively balances quality and computational efficiency by adapting the sampling process to the inherent characteristics of each motion component.
+Table 2: Ablation study on different architectural variants of our bidirectional feature interaction mechanism.
+The results demonstrate the effectiveness of our full model with bidirectional design.Best results in each category are in bold; second best are underlined.
 Model Holistic Expression Gesture
@@
 CoSync ($G\leftrightarrow E$) 312.46 316.97 421.58
-Table 2: Ablation study on different architectural variants of our bidirectional feature interaction mechanism.
-The results demonstrate the effectiveness of our full model with bidirectional design.Best results in each category are in bold; second best are underlined.
+In this section, we conduct comprehensive ablation studies to validate the effectiveness of AsynFusion’s key components and design choices on BEAT Dataset [21].
+Specifically, we examine (1) the impact of different feature interaction strategies in our Dual-branch DiT Architecture, (2) the efficiency of our asynchronous sampling approach compared to synchronized alternatives.
+[…]
+Impact of Different Feature Fusion Methods.
+In our exploration of improving the coordination between facial expressions and body gestures, we explored three distinct feature fusion approaches: Cross Attention (CA), Feature Concatenation (FC), Gated Fusion (GF).
+Our AsynFusion framework implements Cross Attention, leveraging its bidirectional mechanism to achieve fine-grained control over expression-gesture information flow, thereby producing more natural and expressive animations.
+We also examined two alternative methods: Feature Concatenation, which simply concatenates expression and gesture features before feeding them into the next DiT block, and Gated Fusion, which employs learnable weights through a gating mechanism to control the fusion process:
+$F_{fused}=\sigma(W_{g}[F_{exp};F_{ges}])\odot F_{exp}+$ (19)
+$(1-\sigma(W_{g}[F_{exp};F_{ges}]))\odot F_{ges}$
+where $\sigma$ represents the sigmoid function, and $W_{g}$ are the learnable weights.
+As shown in Table 3 This approach adaptively controls each modality’s influence through learnable weights.
+While Gated Fusion achieves better modality balance than simple concatenation, it lacks the sophisticated bidirectional interaction of Cross Attention, making it less capable of capturing the subtle dependencies in natural human behavior.
+Table 3: Ablation study on different Fusion Strategy. Best results in each category are in bold; second best are underlined.
+Fusion Strategy Holistic Expression Gesture
+FMD $\downarrow$ FED $\downarrow$ FGD $\downarrow$
+Feature Concat 340.16 340.23 467.33
+Gated Fusion 325.46 329.15 443.74
+Cross Attention 312.46 316.97 421.58
+Efficiency of Asynchronous LCM Sampling.
+We evaluate three sampling strategies: (1) DDIM (25 steps), (2) Synchronized LCM (8 steps), and (3) our Asynchronous LCM (4 steps for expression, 8 for gesture).
+As shown in Table 4, DDIM achieves the best quality (FMD = 312.46) but is slow (56.4s).
+Synchronized LCM reduces time by 67% (18.6s) with minimal quality drop (FMD = 318.13).
+Our Asynchronous LCM further improves efficiency (15.9s, 72% faster than DDIM) while maintaining competitive quality (FMD = 320.59), showing the benefit of adapting sampling to the convergence speed of each branch.
+Table 4: Comparison of different sampling strategies.
 Sampling Strategy Steps (E/G) Time (s) FMD $\downarrow$
@@
 Async LCM 4/8 15.9 320.59
-Table 3: Comparison of different sampling strategies.
-Method Inference Time (s) GFLOPs
-EMAGE [33] 20.3 18.3
-DiffSHEG [9] 28.6 24.7
-ProbTalk [37] 5.1 9.7
-Our(Sync LCM) 18.6 12.5
-Our(Async LCM) 15.9 10.6
-Table 4: Comparison of computational efficiency.
-Computational Efficiency Comparison.
-Finally, we compare AsynFusion’s computational efficiency against state-of-the-art methods for co-speech motion synthesis.
-As illustrated in Table 4, our method achieves the fastest inference time (15.9s per sequence) and lowest computational complexity (58.4 GFLOPs) while maintaining competitive or superior quality metrics.
-Compared to unified approaches like Probtalk, AsynFusion reduces inference time by 62% and computational requirements by 57%.
-Even against recent efficient methods like DiffSHEG, our approach shows a 43% reduction in inference time and 38% decrease in GFLOPs.
-This significant improvement stems from our dual-branch architecture and asynchronous sampling strategy, which effectively parallelize computation and optimize sampling steps for each motion component.
-These results confirm that AsynFusion successfully addresses the computational efficiency challenges in audio-driven avatar animation while preserving high-quality motion synthesis.
 6 Conclusion
-In this paper, we present AsynFusion, a novel framework for audio-driven avatar animation that advances joint expression-gesture generation.
-Our key contributions include a dual-branch DiT architecture with bidirectional feature interaction, enabled by the CoSync module, which improves the coordination between facial expressions and body gestures.
-Additionally, we introduce an asynchronous sampling strategy using Latent Consistency Models, significantly reducing computational overhead while maintaining high-quality motion synthesis.
+A key limitation of AsynFusion lies in its dependency on training data quality.
+Our model, like other deep learning approaches, inherits both desired behaviors and undesirable artifacts from the training datasets.
+For example, when trained on BEAT [21], the generated motions exhibit jittering artifacts similar to those in the dataset’s female character animations.
+Similarly, expression discontinuities from the SHOW dataset appear in our synthesized results.
+These observations highlight that future improvements in motion synthesis may rely as much on better data collection and cleaning as on model architecture advances.
+Future research could focus on end-to-end multimodal foundation models for motion synthesis, processing speech, text, and video simultaneously.
+Large-scale pretraining across diverse data sources could enable deeper understanding of verbal and non-verbal communication patterns.
+This approach could enhance motion diversity and naturality through universal representations, while advanced cross-modal pretraining could improve the capture of speech-motion correlations for more nuanced animations.
+Acknowledgement
+This work was supported by the National Natural Science Foundation of China under Grant No.
+62441617 and 62476224.
+It was supported by the Postdoctoral Fellowship Program and China Postdoctoral Science Foundation under Grant No.
+2024M764093 and Grant No.
+BX20250485, the Beijing Natural Science Foundation under Grant No.
+4254100, the Fundamental Research Funds for the Central Universities under Grant No.
+KG16336301, and by Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing.